# When Vision Fails: Text Attacks Against ViT and OCR

### v1

@@ Abstract
While text-based machine learning models that operate on visual inputs of rendered text have become robust against a wide range of existing attacks, we show that they are still vulnerable to visual adversarial examples encoded as text. We use the Unicode functionality of combining diacritical marks to manipulate encoded text so that small visual perturbations appear when the text is rendered. We show how a genetic algorithm can be used to generate visual adversarial examples in a black-box setting, and conduct a user study to establish that the model-fooling adversarial examples do not affect human comprehension. We demonstrate the effectiveness of these attacks in the real world by creating adversarial examples against production models published by Facebook, Microsoft, IBM, and Google.
1 Introduction
Advances in adversarial examples [1, 2, 3] for text-based machine learning systems have led to imperceptible perturbation techniques targeting the encoding of text [4, 5]. These attacks leverage uncommon Unicode encodings to subvert text models without visual indication to users [6]. Existing defenses against encoding attacks achieve protection by unifying the visual and encoding pipelines; specifically, the Vision Transformers (ViTs) architecture can be used to build robust new models [7], and Optical Character Recognition (OCR) can be used to retrofit existing models with defenses [4]. Such defenses seek to ensure that text inputs which are visually-identical when rendered produce equivalent outputs in defended models.
@@ 1 Introduction
We show that these defenses are inadequate. We propose a technique to encode text perturbations in a way that, once rendered, the image of the rendered text will contain adversarial perturbations that defeat visual defenses. These adversarial examples operate in the visual text domain, meaning that visual inputs are generated exclusively from rendered text such that it is impossible to perturb arbitrary pixel values. However, by leveraging combining marks from the Unicode specification, we can craft small, targeted visual perturbations that appear on the rendered image of text. While these marks are not sufficiently visually significant to impact a human reader’s understanding of text, the manipulated pixels in the image domain of rendered text enable targeted attacks on model outputs.
@@
Text attacks on ViT and OCR have broad implications. For example, we demonstrate extensions to attacks on toxic content detection [8] and machine translation [9] that succeed despite ViT and OCR defenses against encoding attacks. The techniques we propose can even be used to affect the physical domain such as disclosures submitted within legal cases. If the disclosing legal team adds appropriate text perturbations to printed disclosure documents, the text can form visual adversarial examples that are incorrectly processed by OCR systems despite being readable by humans. This would prevent searchability thus impeding the opposing legal team in the discovery process.
In this paper we make the following contributions:
@@
  • We demonstrate that previously-proposed defenses [4, 7, 10] for encoding attacks against text models including ViT, OCR, and neural encoders are insufficient.
  • We conduct a user study to validate that our adversarial examples do not impact human readability.
@@
2 Background
Figure 1: The visualization gap, against which visual models claim to be robust, continues to be source of adversarial examples via diacritics as in this toxic content classifier example.
The Visualization Gap. Traditionally, machine learning models processing text such as natural language operate directly upon some encoding of the input text. This may take the form of input embeddings as vectors representing words, characters, or learned subword components created by parsing Unicode inputs [6]. However, unlike models, humans do not directly consume encoded text. Rather, text is rendered and then visually presented to human users.
@@ 2 Background
The difference between the encoding and visualization of text can be used to create adversarial examples against models that operate directly upon some form of textual input [4, 5], improving the stealth of earlier techniques leveraging misspelling or paraphrasing [11, 12, 13, 14]. The visualization gap is depicted in Figure 1.
Defense Through Vision. To defend against adversarial examples that exploit the visualization gap, model designers must seek to unify the text encoding and visualization pipelines. That is, designers must seek to build or augment models such that:
$\forall{T}\in t,\ \ E(U(t))=\{t^{\prime}\}\Leftrightarrow\{v(t)\}$ (2)
@@
One simple but effective way to accomplish this on existing models is to render text inputs and process the resulting images through OCR as a pre-processing step prior to model inference [4]. In effect, this provides an automated system that maps fixed visual renderings to a common encoded input. The inference pipeline in this setting is: encoded input $\rightarrow$ rendered image $\rightarrow$ text $\rightarrow$ model.
For greenfield models, Vision Transformers may be the preferred defense approach as no compute-intensive pre-processing model is required. ViTs operate upon images as input, and operating directly upon rendered images as embeddings yields both good performance and defense-by-design against attacks exploiting the visualization gap [7]. The inference pipeline in this setting is: encoded input $\rightarrow$ rendered image $\rightarrow$ model.
Finally, neural encoders offer new NLP models some robustness against Unicode perturbations [10]. Although they do not operate in the visual domain, neural encoders are a form of learned embedding that map relationships between Unicode characters such that encoded values that would look similar if rendered should result in similar embeddings. The inference pipeline in this setting is: encoded input $\rightarrow$ neural embedding $\rightarrow$ model.
Figure 2: Our attack: adversarial examples in the visual domain encoded in the textual domain.
3 Methods
Exploiting Diacritics. In the image domain, adversarial examples are typically crafted by slightly perturbing the values of key pixels often identified through a gradient-based approach [3]. While such an approach would in theory work against ViTs and OCR models, the visual text domain has the added constraint that the input images are generated through rendering text. Therefore, it is not possible to arbitrarily perturb pixel values, as pixel values are not directly encoded; rather, the pixels are generated from encoded text.
@@ 3 Methods
In Unicode, characters commonly used with diacritical marks typically have a dedicated character – or code point – in the Unicode specification. However, the concept of combining diacritical marks also exists. These characters modify the character immediately preceding them to add the specified diacritical mark. The Unicode specification defines 256 such combining characters [15, 16, 17, 18, 19].
@@
Combining diacritical marks provide a method by which arbitrary characters can be subtly perturbed in the image domain. These marks will produce a noticeable visual effect, but as we will later demonstrate through a user study these marks do not affect a human’s ability to read the text. However, just as changing particular pixel values can cause general image classifiers to fail, so too can the perturbed pixel values created by diacritical marks affect ViTs and OCR models. Although the size of the perturbation space is smaller than the more continuous space provided by color pixels in unconstrained images, we will later demonstrate through a series of experiments that this space is sufficiently large to generate adversarial examples for text. This attack is visualized for ViT and OCR pipelines in Figure 2.
@@
[…]
@@
    if round($i$) $\geq 0$ then
     $\bar{x}=$ insert($\bar{x},\mathbf{D}_{d},$round($i$))
@@
[…]
@@
  return  $\textsc{perturb}(\mathbf{P}_{\text{argmax}(\mathbf{\bar{f}})})$
@@
Attack Technique. Visual adversarial examples encoded as text can be generated in a black-box model using a gradient-free optimization technique. We built our technique using differential evolution, a genetic optimization algorithm [20]. Adopting a similar approach previously proposed for imperceptible perturbations [4, 21], we use differential evolution to minimize similarity with the reference output for sequence-to-sequence tasks, and to minimize the output probability of the target class for classification tasks.
The algorithm to generate visual adversarial examples encoded as text is provided as Algorithm 1. In summary, this algorithm takes as input a string to perturb, the target output class or reference output sequence, a visual text model, a perturbation budget representing the maximum number of injected diacritics, and a set of diacritics from which to select. A set of optimization parameters specific to differential evolution are also required by the algorithm [20], although reasonable defaults may be used across repeated invocation. In the case of classification models, this algorithm will return a version of the input string perturbed with diacritics up to the allowed budget that minimize the output probability of the supplied target class. In the case of sequence-to-sequence models – such as the machine translation task – the algorithm minimizes the supplied similarity metric with the supplied reference output.
Since the adversarial example generation algorithm does not rely on gradients, it works in a black-box setting, i.e. the attack does not require access to model weights. Consequently, this algorithm can be used both for local models and on commercial models via APIs as demonstrated in the following section.
4 Evaluation
@@
To examine performance against models, we will first generate adversarial example for an OCR model and measure performance degradation. We will then place this model into a pipeline that renders text input, performs OCR, and then calls the downstream model for both machine translation and toxic content detection and evaluate the performance of adversarial examples. Next, we will craft and measure adversarial examples for a ViT performing machine translation directly on rendered text. Finally, we will demonstrate an extension of these attacks to a neural encoder model performing question answering. A summary of the experimental results is given in Table 1. These results are presented over varying budgets in Figure 4, with additional metric-specific visualizations given in the Appendix.
To examine the effect on human comprehension, we will present a user study in which participants are evaluated for their understanding of the adversarial examples produced in the previous experiments.
@@
All adversarial examples were generated on a cluster of machines each equipped with a Tesla P100 GPU and Intel Xeon Silver 4110 CPU running Ubuntu. We followed Algorithm 1 for example generation, selecting a a population size of 32, 10 iterations, a crossover probability of 0.7, a differential weight dithering from 0.5 to 1, and a varying budget ranging from 0 to 5. For each experiment, we generate 500 adversarial examples for each budget value. Perturbations are generated from the subset of diacritical marks in Unicode supported by the Microsoft Arial Unicode font [22], which is U+300-U+346 and U+360-U+361. All experimental code and results are made available at github.com/nickboucher/diacritics.
Model Baseline Perf Adv. Ex. Perf $\beta=5$ Attack Perf Drop TrOCR [23] 0 12 12 TrOCR-FairSeq [24] 60.9 24.2 60.3% TrOCR-IBM Toxic [25] 87.6 31.8 63.7% ViT FairSeq [7] 57.9 24.5 57.7% CANINE SQuAD [10] 67.6 5.0 92.6% Table 1: Model performance against adversarial examples of diacritical mark budget 5.
Figure 3: Evaluation of diacritic adversarial examples against TrOCR.

### v2

@@ Abstract.
Text-based machine learning models are vulnerable to an emerging class of Unicode-based adversarial examples capable of tricking a model into misreading text with potentially disastrous effects. The primary existing defense against these attacks is to preprocess potentially malicious text inputs using optical character recognition (OCR). In theory, OCR models will ignore any malicious Unicode characters and will extract the visually correct input to be fed to the model. In this work, we show that these visual defenses fail to prevent this type of attack. We use a genetic algorithm to generate visual adversarial examples (i.e., OCR outputs) in a black-box setting, demonstrating a highly effective novel attack that substantially reduces the accuracy of OCR and other visual models. Specifically, we use the Unicode functionality of combining characters (e.g., $\tilde{n}$ which combines the characters $n$ and $\sim$) to manipulate text inputs so that small visual perturbations appear when the text is displayed. We demonstrate the effectiveness of these attacks in the real world by creating adversarial examples against production models published by Meta, Microsoft, IBM, and Google. We additionally conduct a user study to establish that the model-fooling adversarial examples do not affect human comprehension of the text, showing that language models are uniquely vulnerable to this type of text attack.
optical character recognition, vision transformer, adversarial examples, unicode, diacritics
journalyear: 2025copyright: ccconference: Proceedings of the 2025 Workshop on Large AI Systems and Models with Privacy and Security Analysis; October 13–17, 2025; Taipei, Taiwanbooktitle: Proceedings of the 2025 Workshop on Large AI Systems and Models with Privacy and Security Analysis (LAMPS ’25), October 13–17, 2025, Taipei, Taiwandoi: 10.1145/3733800.3763269isbn: 979-8-4007-1896-0/2025/10isbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacy Systems securityccs: Computing methodologies Computer vision
1. Introduction
Adversarial examples are inputs to models that contain subtle perturbations designed to cause models to output undesirable results. They were first identified in the image domain, where small changes in pixel values could cause classifiers to fail (Szegedy et al., 2014; Biggio et al., 2013; Goodfellow et al., 2014). Later advances brought adversarial examples to natural language, where text encoded as Unicode (The Unicode Consortium, 2021f) could have characters inserted or replaced to create adversarial examples in the text domain (Boucher et al., 2022; Pajola and Conti, 2021).
Until now, it has been thought that vision model designed to process text were robust against Unicode adversarial examples. Model architectures such as Vision Transformers (ViTs) (Salesky et al., 2021) and Optical Character Recognition (OCR) pre-processing (Boucher et al., 2022) leverage rendered text as input, suggesting that adversarial perturbations at the text encoding level would not affect these models. We show that this is not the case.
@@ 1. Introduction
In this work, we show that these visual models are themselves vulnerable to Unicode-based attacks. We propose a technique to encode Unicode perturbations in a way that, once rendered, the image of the rendered text will contain adversarial perturbations that defeat visual defenses. By leveraging combining marks from the Unicode specification, we can craft small, targeted visual perturbations that appear on the rendered image of text. While these marks are not visually significant enough to impact a human reader’s understanding of text (as we show in a user study), the manipulated pixels in the image domain of rendered text enable targeted attacks on model outputs. This class of adversarial examples is distinct from general image adversarial examples in that pixel values are not perturbed directly, but rather are modified by perturbing pre-rendered Unicode.
In particular, we build on existing adversarial text perturbation techniques to introduce a new class of Unicode-based adversarial examples that leverages Unicode’s combining diacritical mark functionality. Prior work claiming ViT and OCR are robust against text-based perturbations (Boucher et al., 2022; Salesky et al., 2021) did not consider text containing combining diacritics prior to visual rendering, which we leverage in this work to demonstrate OCR’s vulnerability to Unicode-based attacks. We demonstrate that these barely perceptible marks have a substantial impact on ViT and OCR model performance, causing performance drops of around 60% to as high as 92.6%.
Figure 1. The visualization gap, against which visual models claim to be robust, continues to be a source of adversarial examples via diacritics as in this toxic content classifier example.
@@
Text-based attacks on ViT and OCR have broad implications, particularly in cases involving large-scale, automated text input with minimal human review. For example, we demonstrate extensions to attacks on toxic content detection (Hosseini et al., 2017) and machine translation (Belinkov and Bisk, 2018) that succeed despite ViT and OCR defenses against encoding attacks. The techniques we propose can even be used to affect printed text, such as addresses on envelopes sent through the U.S. Postal Service (which uses OCR to speed up processing (United States Postal Service, 2023), or disclosures submitted as part of a legal case. In the latter example, if the disclosing legal team adds targeted text perturbations to printed disclosure documents, the text can form visual adversarial examples that are incorrectly processed by OCR systems despite being readable by humans. Large-scale document scanning in particular is often automated and performed with little to no human review of the original paper documents—and so it is entirely plausible that a system may scan encoded text (and therefore produce incorrect output) without detection. This could, for instance, prevent searchability, thus impeding the opposing legal team in the discovery process.
In summary, we make the following contributions in this paper:
@@
  • We demonstrate that existing defenses (Boucher et al., 2022; Salesky et al., 2021; Clark et al., 2022) for encoding attacks against text models including ViT, OCR, and neural encoders are insufficient.
  • We conduct a user study to validate that our adversarial examples do not impact human readability and comprehension.
@@
2. Background
2.1. The Visualization Gap
Traditionally, machine learning models processing text such as natural language operate directly upon some encoding of the input text. This may take the form of input embeddings as vectors representing words, characters, or learned subword components created by parsing Unicode inputs (The Unicode Consortium, 2021f). However, unlike models, humans do not directly consume encoded text. Rather, text is rendered and then visually presented to human users.
@@ 2.1. The Visualization Gap
The difference between the encoding and visualization of text can be used to create adversarial examples against models that operate directly upon some form of textual input (Boucher et al., 2022; Pajola and Conti, 2021), improving the stealth of earlier techniques leveraging misspelling or paraphrasing (Gao et al., 2018; Li et al., 2018; Belinkov and Bisk, 2017; Khayrallah and Koehn, 2018). The visualization gap is depicted in Figure˜1.
Figure 2. Our attack: adversarial examples in the visual domain encoded in the textual domain.
2.2. Defense Through Vision
To defend against adversarial examples that exploit the visualization gap, model designers must seek to unify the text encoding and visualization pipelines. That is, designers must seek to build or augment models such that:
(2) $\forall{t}\in T,\ \ E(U(t))=\{t^{\prime}\}\Leftrightarrow\{v(t)\}$
@@ 2.2. Defense Through Vision
One simple but effective way to accomplish this on existing models is to render text inputs and process the resulting images through OCR as a pre-processing step prior to model inference (Boucher et al., 2022). In effect, this provides an automated system that maps fixed visual renderings to a common encoded input. The inference pipeline in this setting is: encoded input $\rightarrow$ rendered image $\rightarrow$ text $\rightarrow$ model.
For greenfield models, Vision Transformers may be the preferred defense approach as no compute-intensive pre-processing model is required. ViTs operate upon images as input, and operating directly upon rendered images as embeddings yields both good performance and defense-by-design against attacks exploiting the visualization gap (Salesky et al., 2021). The inference pipeline in this setting is: encoded input $\rightarrow$ rendered image $\rightarrow$ model.
Finally, neural encoders offer new NLP models some robustness against Unicode perturbations (Clark et al., 2022). Although they do not operate in the visual domain, neural encoders are a form of learned embedding that map relationships between Unicode characters such that encoded values that would look similar if rendered should result in similar embeddings. The inference pipeline in this setting is: encoded input $\rightarrow$ neural embedding $\rightarrow$ model.
2.3. Diacritical Marks
@@
In Unicode, characters commonly used with diacritical marks typically have a dedicated character – or code point – in the Unicode specification. However, the concept of combining diacritical marks also exists. These characters modify the character immediately preceding them to add the specified diacritical mark. The Unicode specification defines 256 such combining characters (The Unicode Consortium, 2021a, b, c, d, e).
3. Related Work
3.1. Text-Based Adversarial Examples
While most adversarial examples proposed in prior work have been image classification attacks (Szegedy et al., 2014; Goodfellow et al., 2014; Athalye et al., 2018), recent work has demonstrated the existence of adversarial examples against text-based ML models. In particular, the accuracy of text classification models decreases on noisy inputs and it is possible to deliberately craft adversarial texts targeting systems reliant on text understanding (Gao et al., 2018; Li et al., 2018; Belinkov and Bisk, 2017; Khayrallah and Koehn, 2018; Alzantot et al., 2018; Zou et al., 2019).
In contrast to prior work presenting visibly perceptible perturbations, Boucher et al. (2022) previously showed that the discrepancy between text encoding and text visualization can be exploited as an attack vector against text-based machine learning models. They proposed optical character recognition (OCR) models as a “catch-all defense” against such attacks, arguing that the visual nature of the input would mitigate differences in visual and text encoding representation that left text-based processing systems vulnerable.
While Boucher et al. (2022) evaluated OCR as a defense against imperceptible perturbations, diacritics present a new class of attacks that are often scarcely more perceptible than invisible perturbations but have a substantial impact on OCR accuracy. Salesky et al. (2021) showed that Arabic diacritics and other minor text modifications substantially degrade the performance of text-based models, but concluded that ViTs achieve “relatively robust” performance against diacritized text. In this paper we demonstrate that both OCR systems and ViTs are indeed vulnerable to the Unicode-based adversarial examples introduced by Boucher et al. (2022), and that such examples significantly impact the performance of these models.
3.2. Attacks on OCR
Most prior work in breaking OCR systems has focused on visual CAPTCHAs (Yan and El Ahmad, 2007; Azad and Jain, 2013), but OCR is commonly used in a variety of preprocessing tasks. Recently, Chen et al. (2020) demonstrated a class of attacks on OCR using adversarial watermarks in a twist on classic adversarial image examples. The commonality of watermarks on inputs to OCR systems makes this attack effectively imperceptible to humans, though this attack relies on non-text image data in inputs. In contrast, in this work we present a class of attacks in which the adversarial examples are encoded directly into text rather than perturbing an image’s background. To the best of our knowledge, ours is the first such attack that manipulates pre-rendered text directly, using the resulting image as an adversarial example.
3.3. Toxic Content Detection
Identifying toxic content online has proven to be one of the most pervasive yet elusive research questions in natural language processing. A rapidly growing area of research, toxic content detection is challenging even when classifying standard Unicode characters in the English language, and toxic content classifiers are generally not robust against adversarial examples (Kurita et al., 2019; Risch and Krestel, 2020). Online users commonly deploy a variety of simple text modifications to bypass platform filters: Kurita et al. (2019) gives the adversarial example of writing “s*ut up” in the place of ‘shut up’. While most adversarial examples against toxic content detection are visually perceptible, Boucher et al. (2022) showed it is also possible to modify the Unicode text encoding of toxic content in a way that is visually imperceptible but fools content detection models.
4. Methods
4.1. Exploiting Diacritics
In the image domain, adversarial examples are typically crafted by slightly perturbing the values of key pixels often identified through a gradient-based approach (Goodfellow et al., 2014). While such an approach would in theory work against ViTs and OCR models, the visual text domain has the added constraint that the input images are generated through rendering text. Therefore, it is not possible to arbitrarily perturb pixel values, as pixel values are not directly encoded; rather, the pixels are generated from encoded text.
@@ 4.1. Exploiting Diacritics
Combining diacritical marks provide a method by which arbitrary characters can be subtly perturbed in the image domain. These marks will produce a noticeable visual effect, but as we will later demonstrate through a user study these marks do not affect a human’s ability to read the text. However, just as changing particular pixel values can cause general image classifiers to fail, so too can the perturbed pixel values created by diacritical marks affect ViTs and OCR models. Although the size of the perturbation space is smaller than the more continuous space provided by color pixels in unconstrained images, we will later demonstrate through a series of experiments that this space is sufficiently large to generate adversarial examples for text. This attack is visualized for ViT and OCR pipelines in Figure˜2.
4.2. Attack Technique
@@
  Optimization Params: similarity metric $\mathcal{S}$, population size $s$, evolution iterations $m$, differential weight $F\in[0,2]$, crossover probability $\mathit{CR}\in[0,1]$
@@
        if round(ii) $\geq 0$ then
           $\bar{x}=$ insert($\bar{x},\mathbf{D}_{d},$round(ii))
@@
    $\mathbf{p_{a}},\mathbf{p_{b}},\mathbf{p_{c}}\xleftarrow{\text{rand}}\mathbf{P}$ s.t. $j\neq a\neq b\neq c$
@@
5  return $\textsc{perturb}(\mathbf{P}_{\text{argmax}(\mathbf{\bar{f}})})$
@@
Visual adversarial examples encoded as text can be generated in a black-box model using a gradient-free optimization technique. We built our technique using differential evolution, a genetic optimization algorithm (Storn and Price, 1997). Adopting a similar approach previously proposed for imperceptible perturbations (Boucher et al., 2022; Shumailov et al., 2021), we use differential evolution to minimize similarity with the reference output for sequence-to-sequence tasks, and to minimize the output probability of the target class for classification tasks.
The algorithm to generate visual adversarial examples encoded as text is provided as Algorithm˜1. In summary, this algorithm takes as input a string to perturb, the target output class or reference output sequence, a visual text model, a perturbation budget representing the maximum number of injected diacritics, and a set of diacritics from which to select. A set of optimization parameters specific to differential evolution are also required by the algorithm (Storn and Price, 1997), although reasonable defaults may be used across repeated invocations. We provide the parameters chosen for our implementation in the following section. In the case of classification models, this algorithm will return a version of the input string perturbed with diacritics up to the allowed budget that minimizes the output probability of the supplied target class. In the case of sequence-to-sequence models – such as the machine translation task – the algorithm minimizes the supplied similarity metric with the supplied reference output.
The threat model for this attack contains an adversary who is able to submit inputs to a model, either via an API or locally. The adversary does not have access to the model’s weights (i.e. a black-box setting) and the target model leverages vision-driven defenses for encoding attacks such as OCR pre-processing or a ViT architecture. The adversary’s goal is to craft an input to the model that will cause an incorrect output.
5. Evaluation
Model
Baseline
Perf
Adv. Ex.
Perf $\beta=5$
Attack
Perf Drop
TrOCR (Li et al., 2021) 0 12 n/a
TrOCR-FairSeq (Ott et al., 2019) 60.9 24.2 60.3%
TrOCR-IBM Toxic (IBM, 2020) 87.6 31.8 63.7%
ViT FairSeq (Salesky et al., 2021) 57.9 24.5 57.7%
CANINE SQuAD (Clark et al., 2022) 67.6 5.0 92.6%
Table 1. Model performance against adversarial examples of diacritical mark budget 5.
@@ 5. Evaluation
To examine the first claim, we will generate adversarial examples for an OCR model and measure performance degradation. We will then place this model into a pipeline that renders text input, performs OCR, and then calls the downstream model for both machine translation and toxic content detection and evaluate the performance of adversarial examples. Next, we will craft and measure adversarial examples for a ViT performing machine translation directly on rendered text. Finally, we will demonstrate an extension of these attacks to a neural encoder model performing question answering. A summary of the experimental results is given in Table˜1. These results are presented over varying budgets in Figure˜3, with additional metric-specific visualizations given in Figures˜5, 5, 7, 7 and 8 in the Appendix. We note that since the concept of visual text adversarial examples is new, there are no direct baseline metrics in the literature against which to compare these results.
@@
All adversarial examples were generated on a cluster of machines each equipped with a Tesla P100 GPU and Intel Xeon Silver 4110 CPU running Ubuntu. We followed Algorithm˜1 for example generation, selecting a a population size of 32, 10 iterations, a crossover probability of 0.7, a differential weight dithering from 0.5 to 1, and a varying budget ranging from 0 to 5. Where applicable, we select chrF (Popović, 2015) as the similarity metric $\mathcal{S}$. For each experiment, we generate 500 adversarial examples for each budget value. Perturbations are generated from the subset of diacritical marks in Unicode supported by the Microsoft Arial Unicode font (Microsoft, 2021), which is U+300-U+346 and U+360-U+361. All datasets and models used are publicly available for research purposes. All experimental code and results are made available for future research at github.com/nickboucher/diacritics.
5.1. TrOCR